How much cash does the firm actually have?

You are the CFO of ABC Corp. You are looking to hire a financial analyst, and you’ve given an assignment to two short-listed candidates. You’ve given the following information to the job applicants:

The company reported net sales of $1,875 million. Assume that there were no noncash sales.
Operating costs (excluding depreciation and amortization) were 65% of the company’s total revenues.
Depreciation and amortization charges were 5% of total sales.
Interest charges were 15% of earnings before interest and taxes (EBIT) with a tax rate of 40%.

You’ve asked the candidates to give you a number that best represents the cash flow situation of the company.

Finn submits a report stating that the firm has $381 million in cash available.

Arty submits a report stating that the firm has $286.875 million in cash.

Based on the information given to them, which applicant has provided a better estimate of the company’s current cash flows?
Answer #1

Calculation of cash available:

Revenues = $1,875 million

Less: Operating costs = 1,875*65% = $1,218.75 million

Depreciation and Amortization = 1,875*5% = $93.75 million

EBIT = $562.5 million

Less: Interest = 562.5*15% = $84.375 million

Earnings before tax = $478.125 million

Less: Tax = $191.25 million

Net Income = $286.875 million

Add: Depreciation and Amortization (non-cash expense) = $93.75 million

Cash available = $380.625 million

i.e. $381 million (approx.)

Hence, Finn has provided a better estimate